OpenStudy (anonymous):

Soybean meal is 12% protein; cornmeal is 6% protein. How many pounds of each should be mixed together in order to get 240-lb mixture that is 9%??? helppp!!!

OpenStudy (anonymous):

X=weight of soy Y=weight of cornmeal X+Y=240 0.12*X+0.06*Y=0.09*240

OpenStudy (anonymous):

80 and 160 ??

OpenStudy (anonymous):

80 soybean and 160 cornmeal?

OpenStudy (anonymous):

I got 120 both which makes sense as 12%+6%=18%. Divided in two components is 9%

OpenStudy (anonymous):

oh